divide rs 25000 between a,b and c so that a gets 7/8 of b and c gets 2/3 of a and b.
Answer by dkppathak(439)

divide rs 25000 between a,b and c so that a gets 7/8 of b and c gets 2/3 of a and b.
Answer
as per given conditions we will get
a=7b/8 (1)
a/b=7/8
as per other condition
c=2(a+b)/3 (2)
a+b+C =25000(3)
by putting the value of (1) and (2) in (3)
7b/8+b+2(a+b)/3=25000
7b/8+b+2(7b/8+b)/3=25000
(7b+8+b)(1+2/3)=25000
(7b+8b)/8(5/3)=25000
(15b/8)x5/3=25000
75b/24=25000
b=25000x24/75
b=8000
now a=7b/8
a=7x8000/8=7000
c=(a+b)x2/3
c=(7000+8000)x2/3
c=15000x2/3=10000
Answer
a=7000
b=8000
c=10000
verification
a+b+c=25000
7000+8000+10000=25000
verified